Triple R Ranch sold 1600 head of their Angus and Dairy cattle. In the sale there were 300 more Angus cows sold then dairy cows. How many Angus and Dairy cows were sold in the sale?
Answer by richwmiller(17219)

a+d=1600
300+d=a
300+d+d=1600
2d=1300
d=650
a=950